OpenStudy (anonymous):

Martin combined different amounts of orange juice, pineapple juice, and strawberry juice to make 1850 milliliters (ml) of fruit punch. The amount of orange juice combined was x ml. The amount of pineapple juice combined was 100 ml more than the amount of orange juice, and the amount of strawberry juice was half the amount of orange juice. This is represented in the equation below. x + (x + 100) + 1/2 x = 1850 What was the difference in the amounts of pineapple juice and strawberry juice Martin combined?

OpenStudy (texaschic101):

x + (x + 100) + 1/2x = 1850 x + x + 1/2x = 1850 - 100 2.5x = 1750 x = 1750/2.5 x = 700 amount of orange juice is 700 ml. amount of strawberry juice is 1/2x = 1/2(700) = 350 ml. amount of pineapple juice is 100 + x = 100 + 700 = 800 totaled = 1850 : 700 + 350 + 800 = 1850 1850 = 1850 (correct) ANSWER : difference in pineapple juice and orange juice is 800 - 700 = 100 ml.

OpenStudy (anonymous):

omg thanks!

OpenStudy (texaschic101):

no problem...thats what I am here for :)